# Asymmetric Duos: Sidekicks Improve Uncertainty

### v1

@@ Abstract
The go-to strategy to apply deep networks in settings where uncertainty informs decisions—ensembling multiple training runs with random initializations—is ill-suited for the extremely large-scale models and practical fine-tuning workflows of today. We introduce a new cost-effective strategy for improving the uncertainty quantification and downstream decisions of a large model (e.g. a fine-tuned ViT-B): coupling it with a less accurate but much smaller “sidekick” (e.g. a fine-tuned ResNet-34) with a fraction of the computational cost. We propose aggregating the predictions of this Asymmetric Duo by simple learned weighted averaging. Surprisingly, despite their inherent asymmetry, the sidekick model almost never harms the performance of the larger model. In fact, across five image classification benchmarks and a variety of model architectures and training schemes (including soups), Asymmetric Duos significantly improve accuracy, uncertainty quantification, and selective classification metrics with only ${\sim}10-20\%$ more computation.
@@ 1 Introduction
Deep ensembles [28] have endured as a defacto method in settings that require uncertainty quantification (UQ) for safety or downstream decision making [4, 30, 8]. When models are trained from scratch with different random seeds, this simple strategy yields strong performance across UQ and decision-making benchmarks [38, 21], provides robustness under distribution shifts [41], and requires no modification to model training methodology.
@@
However, as the modern machine learning paradigm increasingly revolves around larger and larger pre-trained models fine-tuned for specific downstream tasks [11, e.g.], deep ensembles become increasingly ineffective and impractical. First, unlike training models from scratch, fine-tuning from a pre-trained model is much less sensitive to hyperparameters and training randomness. While ensembling multiple fine-tuning runs can yield performance increases, it pales in comparison to the gains obtained when ensembling trained-from-scratch models [18, 48]. Second, and more critically, the growing size of modern models makes ensembling more expensive than ever. Even with moderate-sized models like ResNet-50 [23], training and inference costs were a concern [12, 25, 1]; with large architectures such as ViT-B [9], these costs become even more prohibitive. Deploying a single one of these models can approach the limit of practical resources. Ensembling with double, triple, or larger multiples of the computation may be out of the question.
@@
The smallest deep ensemble requires two models, which with standard approaches needs 2×2$\times$2 × the computation for both training and inference. In this paper, we reduce this cost to a fractional amount (e.g. $10-20\%$) while still obtaining meaningful improvements in accuracy, uncertainty quantification, and downstream decision making. Our Asymmetric Duo framework pairs a base model with a smaller “sidekick” to make better predictions together. For instance, we can pair a larger ViT-B with a smaller ResNet-34, while fine-tuning each separately, to improve over either alone. As the sidekick model requires a fraction of the compute at training and inference time, it only requires minimal computational overhead over the base model, which enables practical use.
@@
Intuitively, an Asymmetric Duo might not help or might even hurt. The smaller model’s accuracy can be much lower, and so one might expect that combining predictions from both models could degrade performance. Yet surprisingly, we find that Asymmetric Duos consistently improve performance, even when the base model and sidekick are extremely imbalanced (i.e. the sidekick is 10×10$\times$10 × smaller and $80\%$ as accurate as the base model). Across datasets and model families, Asymmetric Duos yield up to $10{-}15\%$ improvements in accuracy, uncertainty quantification, and selective classification metrics, while only adding $10{-}20\%$ of the computational cost of the base model.
@@
The Asymmetric Duo framework is straightforward to implement: given a working base model, the user adds a second smaller model (by say simply downloading it from a “model zoo”) and fine-tuning it in the same way as the base model. Moreover, Asymmetric Duos are complementary to other compute-efficient model combinations, like Model Soups [58], and apply broadly by incorporating these methods as base or sidekick models. Our experiments show that Duos can improve the uncertainty quantification of both uncertainty-blind pipelines (like standard training) and uncertainty-aware pipelines (like soups), and so offer a practical drop-in addition to existing models.
@@ 2 Background
(We note that alternative measures of predictive uncertainty, such as the entropy of the categorical distribution defined by $\sigma(f(X))$, are common in other works [28, 2, 55]. We explore entropy uncertainty in Appendix C and find no meaningful difference with respect to our main results.)
@@ 2.1 Deep Ensembles
A deep ensemble [28] refers to $m$ independently-trained deep networks $f_{1},\ldots,f_{m}$, typically of identical architecture and training procedure. At test time, each input $X$ is passed through each network and the resulting logits are averaged to produce an aggregate prediction $\bar{f}(X):=\frac{1}{m}\sum_{i=1}^{m}f_{i}(X)$ from which a class prediction and uncertainty estimate can be derived. The randomness of initialization and shuffling of training samples is usually sufficient to produce diverse logits that, when averaged together, produce more accurate class predictions and better calibrated uncertainty estimates than any of the component models in isolation [28, 41]. Deep ensembles have also been shown to be robust to distribution shifts [41, 2, 31] and have thus been adopted across safety-critical and decision-making applications [4, 30, 8].
As stated in the introduction, deep ensembles are often impractical with the modern transfer learning setup [40, 48, 58]. The primary downside is computational cost. It is common for practitioners to fine-tune large pre-trained models, typically downloaded from a “model zoo”, and training or deploying multiple of these large models can be prohibitively expensive [6]. In addition, member models fine-tuned from a shared pre-trained model tend to settle within the same loss basin, preventing the necessary diversity to produce meaningful improvements [39]. Consequently, deep ensembles are often restricted to settings with moderate-sized models trained from scratch [28, 41, 48].
@@ 2.2 Alternative Methods
Model soups [58] average the parameters of independently fine-tuned models, rather than their predictions, yielding a “weight-space” ensemble with the same inference cost as a single model. This method takes advantage of the fact that models fine-tuned from an identical base model often settle into the same pre-trained loss basin, and thus the loss is approximately locally convex with respect to the parameters. This parameter averaging often yields competitive performance with the prediction averaging of ensembling when the component models are fine-tuned, though the gain is less than ensembling trained-from-scratch models.
Shallow ensembles [29] reduce the multiple model training and inference costs of deep ensembles by instead making multiple predictions from a single model. Specifically, the last layer (or the last $\ell$ layers) of a deep network are replaced with $m$ copies that each receive the same input. The outputs of the $m$ “heads” are averaged together and learned jointly. Since only the last layer is duplicated, the computation and parameters added by a shallow ensemble are only a fraction of that of the original deep network. However, the benefits provided by shallow ensembles are weaker and less consistent than those of deep ensembles. In fact, single model baselines sometimes outperform them on selective classification tasks and robustness under distribution shifts [38].
Other methods. Implicit ensembles aim to produce multiple diverse networks at inference that are all derived from a single training run. These networks are derived through Monte Carlo sampling (e.g. when combined with Dropout [12]), checkpoints from the training process [25], or predictions made at early layers [1]. While these approaches have lower training costs than standard deep ensembles, they retain the same inference costs and typically are incompatible with fine-tuning. Weight-sharing ensembles aim to embed multiple models within a single architecture that can be trained simultaneously [56, 10, 53, e.g.]. While some methods incur lower inference-time costs [22], they typically require from-scratch training and bespoke architectures and/or training techniques. Non-ensemble methods often modify architectures, training, or inference; while they can be effective, these modifications may complicate practical use. See [14, 43] for thorough surveys.
@@ 3.1 Aggregating Predictions from Asymmetric Models
In standard deep ensembles, all models are inherently “equal” in that they only differ in randomization. By contrast, the two networks in an Asymmetric Duo explicitly differ in size and predictive performance. For example, an Asymmetric Duo might combine a ConvNeXt-Base (89 million parameters, 15 billion FLOPs, 84.1% Top-1 accuracy on ImageNet1K) [34] and a RegNeXt-1.6GF (9 million parameters, 1.6 billion FLOPs, 79.7% Top-1 accuracy on ImageNet1K) [45]. While this asymmetry achieves computational efficiency, the larger model is more likely to be correct than the other. Aggregating predictions with the equal-weighted logit average of deep ensembles could therefore be detrimental.
@@
$\hat{Y}_{\text{Duo}}(X)=\operatorname*{arg\,max}_{i}[f_{\text{Duo}}(X)]_{i},\qquad\mathrm{unc}({f}_{\text{Duo}}(X))=1-[\sigma(f_{\text{Duo}}(X)]_{\hat{Y}_{\text{Duo}}(X)}.$
@@
We tune $T_{\text{large}}$ and $T_{\text{small}}$ to minimize the negative log likelihood on the same set used for hyperparameter selection during fine-tuning. As there are only two parameters, the learned weights will likely generalize even from a small reused validation set. With $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ validation outputs saved, this tuning takes only seconds with any standard optimizer. Note that this procedure closely matches temperature scaling for (single-model) calibration [20].
This temperature tuning automatically guards against defective Duos. If $f_{\mathrm{small}}$ is too inaccurate, then the Duo can effectively revert to $f_{\mathrm{large}}$ through the weighting $T_{\mathrm{large}}=1$, $T_{\mathrm{small}}=0$. Conversely, any non-trivial weighting (i.e. $T_{\mathrm{small}}>0$) implies that the Asymmetric Duo provides beneift over $f_{\mathrm{large}}$ since it improves negative log likelihood on the validation set.
@@ 3.2 Measuring the Computational Cost of Asymmetric Duos
The cost of a Duo versus a single model is dictated by the degree of asymmetry in size. How to measure size is non-trivial, because no one property fully captures a deep network’s size or complexity. Parameter count, network depth, floating point operations per forward pass (FLOPs), and throughput each reflect different aspects of model capacity and computational cost [54, 50, 24]. To focus on practical use we measure computation. We profile FLOPs due to their consistency across hardware in Section 4, and we profile throughput as measure of real-world time efficiency in Appendix B. Asymmetric duos deliver efficient improvement for either measure of computational cost.
@@
FLOPS Balance=Computation $\text{FLOPS Balance}=\text{\% Computation Increase}=\mathrm{FLOPs}(f_{\mathrm{small}})/\mathrm{FLOPs}(f_{\mathrm{large}}).$
@@ 4 Experiment Setup
We evaluate Asymmetric Duos in the context of image classification, measuring their effects on accuracy, uncertainty quantification, and downstream decision-making. To demonstrate their “out-of-the-box” nature, we focus on Duos that combine models from popular pre-trained models available from the torchvision [42, 36] and timm [57] “model zoos.”
@@
We benchmark ImageNet [7] for large-scale evaluation, and Caltech 256 [19] and iWildCam [3] for measuring transfer to new domains. As accuracy and uncertainty quantification tend to degrade under distribution shifts [41], we further evaluate with ImageNet V2 [47] and iWildCam-OOD [26] to test robustness under such shifts.
@@
For $f_{\mathrm{large}}$ architecture and parameters, we consider pre-trained models that 1. consistently achieve high performance on fine-tuning benchmarks [e.g. 17] and 2. are commonly used in practice for various computer vision tasks. We therefore experiment with four strong $f_{\mathrm{large}}$ that cover convolution and attention architectures: a ConvNeXt-Base [34] pre-trained on ImageNet, a SwinV2-Base [33] pre-trained on ImageNet, an EfficientNet-V2-L [51] pre-trained on the larger-scale ImageNet21K, and a CLIP ViT-B [44, 9] pre-trained on the larger still LAION-2B [49]. These span standard scale (ImageNet) to extremely large scale (LAION). For $f_{\mathrm{small}}$ models, we consider smaller architectures from torchvision and timm [60, 50, 45, 59, 23, 52, 32] (see Appendix A for a complete list). Within a given model family (e.g. ResNet), we choose models of varying size (e.g. ResNet-34, ResNet-50, ResNet-101, etc.) to produce Duos with various levels of asymmetry.
@@
For our downstream datasets evaluations, we fine-tune our models following the LP-FT recipe [27] with standard unweighted Cross-Entropy Loss, AdamW [35] optimizer, linear warm-up with cosine annealing learning rate scheduler, and perform hyperparameter tuning for all models over learning rate and weight decay on a withheld validation set.
We calibrate all fine-tuned models by temperature scaling [20], minimizing the negative log likelihood on the validation set using L-BFGS, to improve baseline uncertainty quantification.
@@
Figure 2 summarizes the class prediction accuracy of all the pre-trained and fine-tuned models that we use to construct Asymmetric Duos. We report accuracy for ImageNet, ImageNet V2, and Caltech 256 and Macro F1 Score for iWildCam2020-WILDS to account for the significant class imbalances present [26].
Consistent with prior work [17, 37], we find a strong correlation between model size (as measured by FLOPs/parameter counts) and in-distribution (InD) and out-of-distribution (OOD) accuracy. In the next section, we investigate to what degree combining pairs of models into Asymmetric Duos allow us to practically traverse the size/performance tradeoff.

### v2

@@ Abstract
The go-to strategy to apply deep networks in settings where uncertainty informs decisions—ensembling multiple training runs with random initializations—is ill-suited for the extremely large-scale models and practical fine-tuning workflows of today. We introduce a new cost-effective strategy for improving the uncertainty quantification and downstream decisions of a large model (e.g. a fine-tuned ViT-B): coupling it with a less accurate but much smaller “sidekick” (e.g. a fine-tuned ResNet-34) with a fraction of the computational cost. We propose aggregating the predictions of this Asymmetric Duo by simple learned weighted averaging. Surprisingly, despite their inherent asymmetry, the sidekick model almost never harms the performance of the larger model. In fact, across five image classification benchmarks and a variety of model architectures and training schemes (including soups), Asymmetric Duos significantly improve accuracy, uncertainty quantification, and selective classification metrics with only ${\sim}10-20\%$ more computation. Code is available at https://github.com/timgzhou/asymmetric-duos.
@@ 1 Introduction
Deep ensembles [lakshminarayanan2017simple] have endured as a de facto method in settings that require uncertainty quantification (UQ) for safety or downstream decision making [carrete2023deep, li2024autonomous, dolezal2022uncertainty]. When models are trained from scratch with different random seeds, this simple strategy yields strong performance across UQ and decision-making benchmarks [mucsanyi2024benchmarking, gustafsson2020evaluating], provides robustness under distribution shifts [ovadia2019can], and requires no modification to model training methodology.
@@
However, as the modern machine learning paradigm increasingly revolves around larger and larger pre-trained models fine-tuned for specific downstream tasks [fang2023does, e.g.], deep ensembles become increasingly ineffective and impractical. First, unlike training models from scratch, fine-tuning from a pre-trained model is much less sensitive to hyperparameters and training randomness. While ensembling multiple fine-tuning runs can yield performance increases, it pales in comparison to the gains obtained when ensembling trained-from-scratch models [gontijo-lopes2022no, sadrtdinov2023to]. Second, and more critically, the growing size of modern models makes ensembling more expensive than ever. Even with moderate-sized models like ResNet-50 [he2016deep], training and inference costs were a concern [gal2016dropout, huang2017snapshot, antoran2020depth]; with large architectures such as ViT-H [dosovitskiy2021an], these costs become even more prohibitive. Deploying a single one of these models can approach the limit of practical resources. Ensembling with double, triple, or larger multiples of the computation may be out of the question.
@@
The smallest deep ensemble requires two models, which with standard approaches needs $2\times$ the computation for both training and inference. In this paper, we reduce this cost to a fractional amount (e.g. $10-20\%$) while still obtaining meaningful improvements in accuracy, uncertainty quantification, and downstream decision making. Our Asymmetric Duo framework pairs a base model with a smaller “sidekick” to make better predictions together. For instance, we can pair a larger ViT-B with a smaller ResNet-34, while fine-tuning each separately, to improve over either alone. As the sidekick model requires a fraction of the compute at training and inference time, it only requires minimal computational overhead over the base model, which enables practical use.
@@
Intuitively, an Asymmetric Duo might not help or might even hurt. The smaller model’s accuracy can be much lower, and so one might expect that combining predictions from both models could degrade performance. Yet surprisingly, we find that Asymmetric Duos consistently improve performance, even when the base model and sidekick are extremely imbalanced (i.e. the sidekick is $10\times$ smaller and $80\%$ as accurate as the base model). Across datasets and model families, Asymmetric Duos yield up to $10{-}15\%$ improvements in accuracy, uncertainty quantification, and selective classification metrics, while only adding $10{-}20\%$ of the computational cost of the base model.
@@
The Asymmetric Duo framework is straightforward to implement: given a working base model, the user adds a second smaller model (by say simply downloading it from a “model zoo”) and fine-tuning it in the same way as the base model. Moreover, Asymmetric Duos are complementary to other compute-efficient model combinations, like Model Soups [wortsman2022model], and apply broadly by incorporating these methods as base or sidekick models. Our experiments show that Duos can improve the uncertainty quantification of both uncertainty-blind pipelines (like standard training) and uncertainty-aware pipelines (like soups), and so offer a practical drop-in addition to existing models.
@@ 2 Background
(We note that alternative measures of predictive uncertainty, such as the entropy of the categorical distribution defined by $\sigma(f(X))$, are common in other works [lakshminarayanan2017simple, basora2025benchmark, wang2024credal]. We explore entropy uncertainty in Appendix C and find no meaningful difference with respect to our main results.)
@@ 2.1 Deep Ensembles
A deep ensemble [lakshminarayanan2017simple] refers to $m$ independently-trained deep networks $f_{1},\ldots,f_{m}$, typically of identical architecture and training procedure. At test time, each input $X$ is passed through each network and the resulting logits are averaged to produce an aggregate prediction $\bar{f}(X):=\frac{1}{m}\sum_{i=1}^{m}f_{i}(X)$ from which a class prediction and uncertainty estimate can be derived. The randomness of initialization and shuffling of training samples is usually sufficient to produce diverse logits that, when averaged together, produce more accurate class predictions and better calibrated uncertainty estimates than any of the component models in isolation [lakshminarayanan2017simple, ovadia2019can]. Deep ensembles have also been shown to be robust to distribution shifts [ovadia2019can, basora2025benchmark, li2023muben] and have thus been adopted across safety-critical and decision-making applications [carrete2023deep, li2024autonomous, dolezal2022uncertainty].
As stated in the introduction, deep ensembles are often impractical with the modern transfer learning setup [neyshabur2020being, sadrtdinov2023to, wortsman2022model]. The primary downside is computational cost. It is common for practitioners to fine-tune large pre-trained models, typically downloaded from a “model zoo”, and training or deploying multiple of these large models can be prohibitively expensive [dehghani2023scaling]. In addition, member models fine-tuned from a shared pre-trained model tend to settle within the same loss basin, preventing the necessary diversity to produce meaningful improvements [mustafa2020deepensembleslowdatatransfer]. Consequently, deep ensembles are often restricted to settings with moderate-sized models trained from scratch [lakshminarayanan2017simple, ovadia2019can, sadrtdinov2023to].
@@ 2.2 Alternative Methods
Model soups [wortsman2022model] average the parameters of independently fine-tuned models, rather than their predictions, yielding a “weight-space” ensemble with the same inference cost as a single model. This method takes advantage of the fact that models fine-tuned from an identical base model often settle into the same pre-trained loss basin, and thus the loss is approximately locally convex with respect to the parameters. This parameter averaging often yields competitive performance with the prediction averaging of ensembling when the component models are fine-tuned, though the gain is less than ensembling trained-from-scratch models.
Shallow ensembles [lee2015m] reduce the multiple model training and inference costs of deep ensembles by instead making multiple predictions from a single model. Specifically, the last layer (or the last $\ell$ layers) of a deep network are replaced with $m$ copies that each receive the same input. The outputs of the $m$ “heads” are averaged together and learned jointly. Since only the last layer is duplicated, the computation and parameters added by a shallow ensemble are only a fraction of that of the original deep network. However, the benefits provided by shallow ensembles are weaker and less consistent than those of deep ensembles. In fact, single model baselines sometimes outperform them on selective classification tasks and robustness under distribution shifts [mucsanyi2024benchmarking].
Other methods. Implicit ensembles aim to produce multiple diverse networks at inference that are all derived from a single training run. These networks are derived through Monte Carlo sampling (e.g. when combined with Dropout [gal2016dropout]), checkpoints from the training process [huang2017snapshot], or predictions made at early layers [antoran2020depth]. While these approaches have lower training costs than standard deep ensembles, they retain the same inference costs and typically are incompatible with fine-tuning. Weight-sharing ensembles aim to embed multiple models within a single architecture that can be trained simultaneously [wen2020batchensemble, dusenberry2020efficient, tran2022plex, e.g.]. While some methods incur lower inference-time costs [havasi2021training], they typically require from-scratch training and bespoke architectures and/or training techniques. Non-ensemble methods often modify architectures, training, or inference; while they can be effective, these modifications may complicate practical use. See [gawlikowski2023survey, pugnana2024deep] for thorough surveys.
@@ 3.1 Aggregating Predictions from Asymmetric Models
In standard deep ensembles, all models are inherently “equal” in that they only differ in randomization. By contrast, the two networks in an Asymmetric Duo explicitly differ in size and predictive performance. For example, an Asymmetric Duo might combine a ConvNeXt-Base (89 million parameters, 15 billion FLOPs, 84.1% Top-1 accuracy on ImageNet1K) [liu2022convnet] and a RegNeXt-1.6GF (9 million parameters, 1.6 billion FLOPs, 79.7% Top-1 accuracy on ImageNet1K) [radosavovic2020designing]. While this asymmetry achieves computational efficiency, the larger model is more likely to be correct than the other. Aggregating predictions with the equal-weighted logit average of deep ensembles could therefore be detrimental.
@@
$\hat{Y}_{\text{Duo}}(X)=\operatorname*{arg\,max}_{i}[f_{\text{Duo}}(X)]_{i},\qquad\mathrm{unc}({f}_{\text{Duo}}(X))=1-[\sigma(f_{\text{Duo}}(X))]_{\hat{Y}_{\text{Duo}}(X)}.$
@@
We tune $T_{\text{large}}$ and $T_{\text{small}}$ to minimize the negative log likelihood on the same set used for hyperparameter selection during fine-tuning. As there are only two parameters, the learned weights will likely generalize even from a small reused validation set. With $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ validation outputs saved, this tuning takes only seconds with any standard optimizer. Note that this procedure closely matches temperature scaling for (single-model) calibration [guo2017calibration].
This temperature tuning automatically guards against defective Duos. If $f_{\mathrm{small}}$ is too inaccurate, then the Duo can effectively revert to $f_{\mathrm{large}}$ through the weighting $T_{\mathrm{large}}=1$, $T_{\mathrm{small}}=0$. Conversely, any non-trivial weighting (i.e. $T_{\mathrm{small}}>0$) implies that the Asymmetric Duo provides benefit over $f_{\mathrm{large}}$ since it improves negative log likelihood on the validation set.
@@ 3.2 Measuring the Computational Cost of Asymmetric Duos
The cost of a Duo versus a single model is dictated by the degree of asymmetry in size. How to measure size is non-trivial, because no one property fully captures a deep network’s size or complexity. Parameter count, network depth, floating point operations per forward pass (FLOPs), and throughput each reflect different aspects of model capacity and computational cost [vasu2023mobileone, tan2019efficientnet, howard2017mobilenets]. To focus on practical use we measure computation. We profile FLOPs due to their consistency across hardware in Section 4, and we profile throughput as measure of real-world time efficiency in Appendix B. Asymmetric duos deliver efficient improvement for either measure of computational cost.
@@
$\text{FLOPs Balance}=\text{\% Computation Increase}=\mathrm{FLOPs}(f_{\mathrm{small}})/\mathrm{FLOPs}(f_{\mathrm{large}}).$
@@ 4 Experiment Setup
We evaluate Asymmetric Duos in the context of image classification, measuring their effects on accuracy, uncertainty quantification, and downstream decision-making. To demonstrate their “out-of-the-box” nature, we focus on Duos that combine models from popular pre-trained models available from the torchvision [paszke2019pytorch, torchvision2016] and timm [rw2019timm] “model zoos.”
@@
We benchmark ImageNet [deng2009imagenet] for large-scale evaluation, and Caltech 256 [griffin2007caltech] and iWildCam [beery2020iwildcam2020competitiondataset] for measuring transfer to new domains. As accuracy and uncertainty quantification tend to degrade under distribution shifts [ovadia2019can], we further evaluate with ImageNet V2 [recht2019imagenet] and iWildCam-OOD [koh2021wilds] to test robustness under such shifts.
@@
For $f_{\mathrm{large}}$ architecture and parameters, we consider pre-trained models that 1. consistently achieve high performance on fine-tuning benchmarks [e.g. goldblum2023battle] and 2. are commonly used in practice for various computer vision tasks. We therefore experiment with four strong $f_{\mathrm{large}}$ that cover convolution and attention architectures: a ConvNeXt-Base [liu2022convnet] pre-trained on ImageNet, a SwinV2-Base [liu2022swin] pre-trained on ImageNet, an EfficientNet-V2-L [tan2021efficientnetv2] pre-trained on the larger-scale ImageNet21K, and a CLIP ViT-B [radford2021learning, dosovitskiy2021an] pre-trained on the larger still LAION-2B [schuhmann2022laionb]. These span standard scale (ImageNet) to extremely large scale (LAION). For $f_{\mathrm{small}}$ models, we consider smaller architectures from torchvision and timm [zhang2018shufflenet, tan2019efficientnet, radosavovic2020designing, xie2017aggregated, he2016deep, tan2019mnasnet, liu2021swin] (see Appendix A for a complete list). Within a given model family (e.g. ResNet), we choose models of varying size (e.g. ResNet-34, ResNet-50, ResNet-101, etc.) to produce Duos with various levels of asymmetry.
@@
For our downstream datasets evaluations, we fine-tune our models following the LP-FT recipe [kumar2022fine] with standard unweighted Cross-Entropy Loss, AdamW [loshchilov2018decoupled] optimizer, linear warm-up with cosine annealing learning rate scheduler, and perform hyperparameter tuning for all models over learning rate and weight decay on a withheld validation set.
We calibrate all fine-tuned models by temperature scaling [guo2017calibration], minimizing the negative log likelihood on the validation set using L-BFGS, to improve baseline uncertainty quantification.
@@
Figure 2 summarizes the class prediction accuracy of all the pre-trained and fine-tuned models that we use to construct Asymmetric Duos. We report accuracy for ImageNet, ImageNet V2, and Caltech 256 and Macro F1 Score for iWildCam2020-WILDS to account for the significant class imbalances present [koh2021wilds].
Consistent with prior work [goldblum2023battle, miller2021accuracy], we find a strong correlation between model size (as measured by FLOPs/parameter counts) and in-distribution (InD) and out-of-distribution (OOD) accuracy. In the next section, we investigate to what degree combining pairs of models into Asymmetric Duos allow us to practically traverse the size/performance tradeoff.